On January 1, Year 1, Zero Company obtained a $52,000, 4-year, 6.5% installment note from Regional Bank. The note requires annual payments of $15,179, beginning on December 31, Year 1. The December 31, Year 3 carrying amount in the allocation of periodic payments table for this installment note will be equal to

$6,463

$13,000

$14,252

$0
Answer #1

$14,252

Year Beginning notes Payable Interest Expense Annual payment Reduction of notes payable Ending Notes Payable
a b=a*6.5% c d=c-b e=a-d
1 $       52,000 $       3,380 $       15,179 $       11,799 $       40,201
2 $       40,201 $       2,613 $       15,179 $       12,566 $       27,635
3 $       27,635 $       1,796 $       15,179 $       13,383 $       14,252
4 $       14,252 $           926 $       15,179 $       14,253 $                -0